OpenStudy (anonymous):

Suppose that f(5) = 1, f '(5) = 4, g(5) = −9, and g'(5) = 6. Find the following values. (f/g)'(5)=? (g/f)'(5)=?

OpenStudy (anonymous):

would the first one be -8/9 and the next one 72?

OpenStudy (anonymous):

it said they were wrong

OpenStudy (freckles):

that isn't the quotient rule

OpenStudy (freckles):

\[(\frac{f}{g})'=\frac{g \cdot f' -f \cdot g'}{g^2}\]

OpenStudy (anonymous):

and that would be -9(4)-1(6)/(-9^2)

OpenStudy (welshfella):

ah yes - freckles has it

OpenStudy (freckles):

\[(\frac{f}{g})'(5)=\frac{4(-9)-6(1)}{(-9)^2}\] is correct

OpenStudy (anonymous):

the answer is-8/9, but i get it wrong

OpenStudy (freckles):

then answer for (f/g)'(5) isn't -8/9

OpenStudy (phi):

how did you get -8/9 ?

OpenStudy (anonymous):

really? -36-6/81=-72/81

OpenStudy (freckles):

-36-6=-42

OpenStudy (anonymous):

haha wow i, dumb

OpenStudy (anonymous):

i kept thinking 42 but i wrote 72 each time, ok thanks for clearing that up

OpenStudy (phi):

that simplifies to -14/27

OpenStudy (anonymous):

yea got it

OpenStudy (freckles):

I but that is exactly what happened on your second problem too you wrote 72 instead of 42 :p